Exercise 3-20 #11

Reclassification of long-term notes payable to current notes payable

  1. What effect on current ratio?
  2. What effect on quick ratio?
  3. What effect on debt to equity ratio?
Answer #1

Answer

Effects of Reclassification of long term note payable to current note payable

1 effects on current ratio

If the long term note payable is reclassified as current note payable, it leads to increase in current liabilities. Increase in current liabilities DECREASE the current ratio.

Current ratio = current assets / current liabilities

For example

If current assets is $255,678 current liabilities is 67,987

Current ratio = current assets / current liabilities

Current ratio = 255,678 / 67,987 = 3.76 times

If a long term note payable $20,000 is reclassified as current liability. Therefore current liabilities increased to $87,987 (67,987 + 20,000)

Current ratio = 255,678 / 87,987 = 2.90 times

The above example clearly show that reclassification of long term note payable to current note payable is lead to decrease the current ratio.

2 Reclassification of long term note payable to current note payable is leads to increase current liabilities and DECREASE the quick ratio.

Quick ratio = ( cash and cash equivalents + marketable securities + account receivables) / current liabilities

For example

The value of cash and cash equivalents, marketable securities and account receivables is $24,675 and current liabilities is $12,564

Quick ratio = 24,675 / 12,564 = 1.96 times

If long term note payable $5,400 is reclassified as current note payable , the current liabilities is increased to $17,964 (12,564 + 5,400).

Quick ratio = 24,675 / 17,964 = 1.37 times

The above example show that reclassification of long term note payable to current note payable decrease the quick ratio.

3 Reclassification of long term note payable to current note payable NOT EFFECTS the debt to equity ratio.

Debt to equity ratio = total debt / equity

*Total debt includes both current liabilities and long term liabilities.

For example

If long term note payable is $10,000 , current liabilities is 23,000 and other non current liabilities is $15,000.

Equity is $120,000

Total debt ( total liabilities)

= 10,000 + 23,000 + 15,000 = $48,000

Debt to equity ratio = 48,000 / 120,000 = 0.4

It long-term note payable $10,000 is reclassified as current liability, it will increase the current liabilities to $33,000 but not change the total liabilities (total debt).

Total liabilities (total debt) remain at $48,000.

Therefore debt to equity ratio remains 0.4.

In short the reclassification of long term note payable to current note payable decrease the current ratio and quick ratio, and not effect the debt to equity ratio

If YOU are like the answer please upvote.